OpenStudy (anonymous):

1. Renting a bicycle at a local Bike shop costs $30 for the first day and $10 for each additional day. a. Create a chart of values and write an equation to find the total cost, C in terms of the number of days d. b. Use the equation you created in Part a to find the cost of renting a bike for 5 days. SHOW ALL THE WORK PLEASE :)

OpenStudy (igreen):

Equation: C = 10d + 30

OpenStudy (imstuck):

The equation would be written with the $30 as a constant. That means that no matter how many days you have the bike, you will pay 30. How many days is the independent variable. y = 30 + (how much per day)(the cost per day).

OpenStudy (anonymous):

C=10d + 30

OpenStudy (anonymous):

thats the equation correct

OpenStudy (imstuck):

yes that is the equation.

OpenStudy (anonymous):

okay the value chart?

OpenStudy (imstuck):

now you plug in your number of days. Just pick some. Start with one day, which will be your d, and then do the math. y = 30 + 10(1) puts the y at 40. Now do 2 days. y=30 + 10(2) which puts y at 50. Do the same for several days.

OpenStudy (imstuck):



OpenStudy (anonymous):

in 5 days

OpenStudy (anonymous):

so it would be 80 in 5 days

OpenStudy (imstuck):

They want you to plug in 5 for d. Do this: Cost = 30 + 10(5). What is the cost for this equation?

OpenStudy (imstuck):

yes 80

OpenStudy (anonymous):

so thats the value chart up above

OpenStudy (imstuck):

yes it is. Do the 4 and the 5. Put them in the d column, then fill in the appropriate cost (y).

OpenStudy (anonymous):

thank you both so so much :)

OpenStudy (igreen):

So with each passing day, the price increases by 10.